OpenStudy (anonymous):

Find < MNO 30o 60o 90o 120o

OpenStudy (anonymous):

@GodlyMaths

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@Godlovesme

OpenStudy (dumbcow):

inscribed angles equal half the opposite arc

OpenStudy (anonymous):

How do i solve it ?

OpenStudy (dumbcow):

what is the measure of the arc?

OpenStudy (anonymous):

how do i find the measure of the arc?

OpenStudy (godlovesme):

360-120= 240 240-120= 120/2= 60 there u go :) hope that helped

OpenStudy (dumbcow):

it gives it to you

OpenStudy (anonymous):

120

OpenStudy (anonymous):

Thank you @Godlovesme

OpenStudy (dumbcow):

take half of that and you get the inscribed angle

OpenStudy (anonymous):

60?

OpenStudy (dumbcow):

yes

OpenStudy (godlovesme):

My pleasure and u r correct!! C: thanks for the medal

OpenStudy (anonymous):

Can you help me with more please both of yall ? I need it for graudation

OpenStudy (godlovesme):

Sure I'll try

OpenStudy (anonymous):

Find m< 1

OpenStudy (anonymous):

45o 30o 75o 90o

OpenStudy (anonymous):

@Godlovesme

OpenStudy (dumbcow):

dont they give you a geometry reference sheet with all the formulas

OpenStudy (anonymous):

Nope :(

OpenStudy (dumbcow):

ok hold on, ill try to find one for you ...

OpenStudy (anonymous):

Okay

OpenStudy (godlovesme):

Add bother 110 and 40 then divide by 2 :)

OpenStudy (godlovesme):

Both** >.<

OpenStudy (anonymous):

Omg you are awesome

OpenStudy (anonymous):

Find <m

OpenStudy (godlovesme):

Thanks :D and anytime

OpenStudy (anonymous):

18o 28o 38o 8o

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

Hang on

OpenStudy (anonymous):

Okay :)))

OpenStudy (godlovesme):

Sorry can't figure this out :( I apologize @euniceOB97

OpenStudy (anonymous):

Its okay :( @Godlovesme

OpenStudy (anonymous):

Can you please help @dumbcow

OpenStudy (dumbcow):

because LK and LM are both tangent to circle the lengths are equal --> 4x -2 = 3x+3

OpenStudy (anonymous):

Do i substract ?

OpenStudy (anonymous):

When i did that equation i got 5?

OpenStudy (dumbcow):

that is correct, x=5

OpenStudy (anonymous):

What i do next ?

OpenStudy (dumbcow):

i dunno what are they asking for?

OpenStudy (anonymous):

find <m?

OpenStudy (dumbcow):

M is not an angle but a point on the circle?

OpenStudy (anonymous):

I think so

OpenStudy (dumbcow):

they prob want length of LM which is 3x+3 ..... plug in x=5

OpenStudy (anonymous):

3x+3=5?

OpenStudy (dumbcow):

no LM = 3x+3 x=5 plug in 5 for x

OpenStudy (anonymous):

18?

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

True or False. When finding a point equidistant from three noncolinear points, you need to find where the perpendicular bisectors of the segments connecting them intersect.

OpenStudy (anonymous):

Please answer this

OpenStudy (anonymous):

You deserve medals @dumbcow

OpenStudy (anonymous):

Please help @Godlovesme :)

OpenStudy (godlovesme):

@cosmichaotic help please :)

OpenStudy (godlovesme):

Srry idk :(

OpenStudy (dumbcow):

i think its true but not completely sure

OpenStudy (anonymous):

Six posts are evenly spaced along a circular wall encasing a fountain. Anna stands at point A, where she cannot see beyond the only two posts visible to her. What is m<A?

OpenStudy (anonymous):

100o 120o 125o 150o

OpenStudy (dumbcow):

@EuniceOB97 , prob better if you just open a new post for new questions

OpenStudy (anonymous):

Okay great idea i will